\begin{thebibliography}{998}
	
	
%BBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBB

\bibitem{BD2011}Bazzanella, D. \textit{\color{red}Some conditional results on primes between consecutive squares.}
Funct. Approx. Comment. Math. 45 (2011), 255--263. \href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2895157}{MR2895157}.

\bibitem{BH2000} Bays, C.; Hudson, R. \textit{\color{red}A new bound for the smallest $x$ with $\pi(x)>\li(x)$.} Math. Comp. 69 (2000) 1285--1296.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1752093}{MR1752093}.

%\bibitem{BD2000}DANILO BAZZANELLAPrimes between consecutive squares
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR3584567}{MR3584567}.

\bibitem{BP2001} Baker, R. C.; Harman, G.; Pintz, J. \textit{\color{red}The difference between consecutive primes. II.}
Proc. London Math. Soc. (3) 83 (2001), no. 3, 532--562.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1851081}{MR1851081}.

\bibitem{BR2018} Bennett, M.; Martin, G.; O'Bryant, K.; Rechnitzer, A. \textit{\color{red}Explicit bounds for primes in arithmetic progressions.} Illinois J. Math. 62 Number 1--4 (2018) pp 427--532. http://arxiv.org/abs/1802.00085.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR3922423}{MR3922423}.


%CCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCCC

\bibitem{CJ2023} Cully-Hugill, Michaela;  Johnston, D. \textit{\color{red}
	On the error term in the explicit formula of Riemann-von Mangoldt II.}
http://arxiv.org/abs/2402.04272.

%DDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDD

\bibitem{DA2016} Dudek, A. \textit{\color{red}An Explicit Result for Primes Between Cubes.} Functiones et Approximatio Commentarii Mathematici Vol. 55, Issue 2 (Dec 2016), pp. 177--197. http://arxiv.org/abs/1401.4233. 
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR3584567}{MR3584567}.


%\bibitem{DH2000} Davenport, H. \textit{\color{red}Multiplicative number theory.}
%Grad. Texts in Math., 74 Springer-Verlag, New York, 2000
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR1790423}{MR1790423}.

%\bibitem{DL2012}De Koninck, J.; Luca, F. \textit{\color{red}Analytic number theory.} Grad. Stud. Math., 134
%American Mathematical Society, Providence, RI, 2012.
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2919246}{MR2919246}.

\bibitem{DP2016} Dusart, P. \textit{\color{red}Estimates of some functions over primes, without R.H..} Math. Comp. 85 (2016), no. 298, 875--888. http://arxiv.org/abs/1002.0442.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR3434886}{MR3434886}.



%FFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFF
	

	
%EEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEE

\bibitem{EE1985}Ellison, W.; Ellison, Fern. \textit{\color{red}Prime numbers.} A Wiley-Interscience Publication. John Wiley \& Sons, 
Inc., New York; Hermann, Paris, 1985.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0814687}{MR0814687}.

%GGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGGG

%\bibitem{GE1967} Grosswald, E. \textit{\color{red}Oscillation Theorems of Arithmetical Functions}. Transactions of the American Mathematical Society, Vol. 126, No. 1 (Jan., 1967), pp. 1--28.
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0202685}{MR0202685}.

\bibitem{GL2023}Granville, A.; Lumley, Allysa. \textit{\color{red}Primes in short intervals: heuristics and calculations.} Exp. Math. 32 (2023), no. 2, 378--404. http://arxiv.org/abs/2009.05000.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR4592954}{MR4592954}.

%\bibitem{GM1986}???	D. A. GOLDSTON and M. L. MONTGOMERY, Pair correlation of zero and primes in short inter
	


	
%HHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH



%\bibitem{HG2007} Harman, G. \textit{\color{red}Prime-detecting sieves.} London Math. Soc. Monogr. Ser., 33 Princeton University Press, Princeton, NJ, 2007
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2331072}{MR2331072}.
	
%IIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIIII
	
\bibitem{IA1937} Ingham, A. \textit{\color{red}On the difference between consecutive primes}. The Quarterly Journal of Mathematics, (1):255--266, 1937.
\href{https://zbmath.org/0017.38904}{Zbl 0017.38904}.

\bibitem{IA1985} Ivic, A. \textit{\color{red}The Riemann Zeta-Function: Theory and Applications.}
Wiley-Intersci. Publ., New York, 1985. 
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0792089}{MR0792089}.

%\bibitem{IK2004} Iwaniec, H.; Kowalski, E. \textit{\color{red}Analytic number theory.} Amer. Math. Soc. Colloq. Publ., 53 American Mathematical Society, Providence, RI, 2004.
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2061214}{MR2061214}.

%\bibitem{KT2008} Kotnik, T. \textit{\color{red}The prime-counting function and its analytic approximations: $\pi(x)$ and its approximations.} Preprint, Adv. Comput. Math. 29 (2008), no. 1, 55--70.
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2420864}{MR2420864}.

\bibitem{KW2009} Kaczorowski, J.; Wiertelak, K. \textit{\color{red}Oscillations of a given size of some arithmetic error terms.} Trans. Amer. Math. Soc. 361 (2009), no. 9, 5023--5039.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2506435}{MR2506435}.

%LLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLL


%MMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMMM
%\bibitem{MO2019} Mathoverflow. \textit{\color{red}Prime Divisors of Values of $\prod_{n\leq x}\left(n^2+1 \right)$.} 
%\href{https://mathoverflow.net/questions/337291}{Mathoverflow}.




%\bibitem{MO2018} Mathoverflow. \textit{\color{red}The inequality $\pi((n+1)^2)-\pi(n^2) \le \pi(n)$.} 
%\href{https://mathoverflow.net/questions/303635}{Mathoverflow}.

\bibitem{MS2007} Maier, H.; Stewart, C. \textit{\color{red}On intervals with few prime numbers.} J. Reine Angew. Math.
608 (2007), 183--199.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2339473}{MR2339473}.

\bibitem{MT2024} Matomäki, Kaisa; Merikoski, Jori, Teräväinen Joni. \textit{\color{red}Primes in arithmetic progressions and short intervals without $L$-functions.} http://arxiv.org/abs/2401.17570.


\bibitem{MV2007} Montgomery, H.; Vaughan, R. \textit{\color{red}Multiplicative number theory. I. Classical theory}. Cambridge University Press, Cambridge, 2007.
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2378655}{MR2378655}.
%NNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNNN

%RRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRRR

\bibitem{RH1987} te Riele, Herman J. \textit{\color{red}On the sign of the difference $\pi(x)-{\rm li}(x)$.} Math. Comp. 48 (1987), no. 177, 323--328.


%\bibitem{RS1962} Rosser, J. B.; Schoenfeld, L. \textit{\color{red}Approximate formulas for some functions of prime numbers.} Illinois J. Math. 6 (1962), 64--94.	
%\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0137689}{MR0137689}.

%SSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSSS

\bibitem{SA1943} Selberg, A. \textit{\color{red}On the normal density of primes in small intervals, and the difference between consecutive primes.} Arch. Math. Naturvid. 47 (1943), no. 6, 87--105.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0012624}{MR0012624}.

\bibitem{SD2010}Saouter, Y.; Demichel, P. \textit{\color{red}A sharp region where $\pi (x)-{\mathrm {li}}(x)$ is positive}. Math. Comp. April, 2010.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR2684372}{MR2684372}.

\bibitem{SL1976} Schoenfeld, L. \textit{\color{red}Sharper bounds for Chebyshev functions $\theta(x)$ and $\psi(x)$ II.} Math. Comp. 30 (1976), 337--360.
\href{https://mathscinet.ams.org/mathscinet-getitem?mr=MR0457374}{MR0457374}.

\bibitem{SW2024}Sorenson, J.; Webster, J. \textit{\color{red}An algorithm and computation to verify Legendre's Conjecture up to $3.33\times10^{13}$.}
http://arxiv.org/abs/2401.13753.

%TTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTTT

	
\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[top=3.5cm, bottom=3.5cm, left=3cm, right=3cm]{geometry}
\usepackage[latin1]{inputenc}
\usepackage{csquotes}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphics}
\usepackage{float}
\usepackage[hang,flushmargin]{footmisc} 

\usepackage{amsmath,amsthm,amssymb,graphicx, multicol, array}
\usepackage{enumerate}
\usepackage{enumitem}
\usepackage{xcolor}

\usepackage[pagebackref,bookmarks,colorlinks,breaklinks,linktoc=page]{hyperref}
\hypersetup{linkcolor=blue,citecolor=red,filecolor=blue,urlcolor=blue} 

\usepackage{tabto}

\numberwithin{equation}{section}

\usepackage{tikz}
\usepackage{pgfplots}

\usetikzlibrary{matrix}

\usepackage{mathrsfs}

\DeclareMathOperator{\li}{li}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}{Lemma}[section]
\newtheorem{theorem}{Theorem}[section]

\usepackage{fancyhdr}
\usepackage{tocloft}

\title{Legendre Conjecture over Arithmetic Progressions}
\date{}
\author{N. A. Carella}

\begin{document}
\maketitle

\textbf{\textit{Abstract}:} 
Let $1\leq a<q$ be a pair of small integers such that $\gcd(a,q)=1$ and let $x>1$ be a large number. This note discusses the existence of a short sequence of primes $p\equiv a\bmod q$ between two squares $x^2$ and $(x+1)^2$. 
\let\thefootnote\relax\footnote{ \today \date{} \\
\textit{AMS MSC2020}: Primary 11N05; Secondary 11N32 \\
\textit{Keywords}: Prime in short interval; Prime in arithmetic progression; Distribution of prime; Legendre conjecture.}


\section{Introduction }\label{S3377A}
A recent result on the theory of primes in short intervals claims that each short interval $[n^2,(n+1)^2]$ contains the expected number of primes $\pi((n+1)^2)-\pi(n^2)=(1+o(1))x/\log x$ with at most $O(x^{1/5+\varepsilon})$ exceptional values $n\leq x$, where  $x$ is a large number and $\varepsilon>0$, see \cite[Theorem 1.1]{BD2011}. This result proves the Legendre conjecture on the existence of primes between two squares for almost all integers $n\geq1$.\\

This note provides a partial result for an extended form of the Legendre conjecture on arithmetic progression, this contribution is proved in \hyperlink{thm3377D.050}{Theorem} \ref{thm3377D.050} unconditionally. The preliminary results in the section provides explicit inequalities for the prime number theorem in arithmetic progressions.

\section{Primes in Arithmetic Progressions} \label{S3377D} \hypertarget{S3377D}
Let $x\geq1$ be a large number and let $1\leq a<q$ be a pair of relatively prime integers. The prime counting function over arithmetic progressions and the associated weighted prime counting function are defined by
\begin{equation}\label{eq3377D.100b}
	\pi(x,q,a)=\sum_{\substack{p\leq x\\p\equiv a\bmod q}}1. 
\end{equation}
The associated weighted prime counting functions are defined by
\begin{equation}\label{eq3377D.100d}
	\psi(x,q,a)=\sum_{\substack{p^k\leq x\\p\equiv a\bmod q}}\log p\quad \text{ and } \quad\theta(x,q,a)=\sum_{\substack{p\leq x\\p\equiv a\bmod q}}\log p, 
\end{equation}
where $k\geq0$. The logarithm integral is the defined by the expression 
\begin{equation}\label{eq3377D.100f}
	\li(x)=\int_2^x\frac{1}{\log t}\;dt. 
\end{equation}{\tiny }
\begin{theorem} \label{thm3377D.100}\hypertarget{thm3377D.100} Let $a$ and $q$ be integers with $1\leq q \leq 10^5$ and $\gcd(a, q) = 1$. If $x \geq 10^3$, then
	\begin{enumerate}[font=\normalfont, label=(\roman*)]
		\item $\displaystyle \left| \psi(x,q,a)-\frac{x}{\varphi(q)}\right|<0.19  \frac{x}{\log x},$\\[.1cm]
		\item $\displaystyle \left| \theta(x,q,a)-\frac{x}{\varphi(q)}\right|<0.40  \frac{x}{\log x},$\\[.1cm]
		\item $\displaystyle \left| \pi(x,q,a)-\frac{\li(x)}{\varphi(q)}\right|<0.53  \frac{x}{(\log x)^2}.$\\[.1cm]
	\end{enumerate}
\end{theorem}	
\begin{proof}[\textbf{Proof}] This is Corollary 1.7 in \cite{BR2018}. Other cases of the theorem for other ranges of $q$ are also derived in this paper.
\end{proof}

Earlier and similar explicit estimates for the prime counting functions and related topics were developed in \cite[Theorem 5.2]{DP2016}, et alii. Lastly, there are the conditional explicit estimates as in \cite[Theorem 10]{SL1976}.\\

The analysis of the oscillations of the differences 
\begin{equation}\label{eq3377D.150c}
	\psi(x)-x  \quad \text{ and }\quad  \psi(x+z)-\psi(x)
\end{equation}
and other associated differences are difficult analysis and are topics of current research, there is a vast literature on these topic, confer \cite[p.\; 191]{EE1985}, \cite[Theorem 15.11]{MV2007}, \cite[p.\; 306]{IA1985}, \cite{RH1987}, \cite{BH2000}, \cite{KW2009}, \cite{SD2010} and 
other sources for details.  However, the analysis of the differences

\begin{equation}\label{eq3377D.150d}
(x+z)-x \quad \text{ and }\quad  \frac{x+z}{\log(x+z)}-\frac{x}{\log x}
\end{equation} are much simpler since these are strictly monotonic increasing functions. The explicit formula in \hyperlink{thm3377D.100}{Theorem} \ref{thm3377D.100} facilitate the derivations of both lower bound and upper bound of the complicated difference $\psi(x+z)-\psi(x)$ in terms of the simpler strictly monotonic functions as demonstrated here.
\begin{lem} \label{lem3377D.150A} \hypertarget{lem3377D.150A} Let $a$ and $q$ be integers with $1\leq q \leq 10^5$ and $\gcd(a, q) = 1$, and let $\varepsilon>0$ be a small number. If $x$ is a large number and $x^{\varepsilon}\leq z\leq x$, then
	\begin{enumerate}[font=\normalfont, label=(\roman*)]
	\item $\displaystyle \psi(x+z,q,a)-\psi(x,q,a)\geq  \frac{x+z}{\varphi(q)}-\frac{x}{\varphi(q)} -(0.19)\left( \frac{x+z}{\log (x+z)}-\frac{x}{\log x}\right),$\\[.1cm]
		\item $\displaystyle \psi(x+z,q,a)-\psi(x,q,a)\leq  \frac{x+z}{\varphi(q)}-\frac{x}{\varphi(q)} +(0.19)\left( \frac{x+z}{\log (x+z)}-\frac{x}{\log x}\right).$
\end{enumerate}
\end{lem}

\begin{proof}[\textbf{Proof}] The functions $f(x)=x/\varphi(q)\geq0$ and $g(x)=x/\log x\geq0$ are strictly monotonically increasing for all sufficiently large real numbers $x\geq x_0(q)$. Moreover, the sums and differences
	\begin{equation}\label{eq3377D.150h}
		f(x+z)-f(x)>0, \quad g(x+z)-g(x)>0, \quad f(x)\pm g(x)>0
	\end{equation}		
	are strictly monotonically increasing nonnegative quantities for all sufficiently large real numbers $x\geq x_0(q)$. The claims in \eqref{eq3377D.150h} are routine exercises. \\
	
	Taking the difference of the inequalities
	
	\begin{equation}\label{eq3377D.150i}
		\frac{x+z}{\varphi(q)}-0.19  \frac{x+z}{\log (x+z)}\leq\psi(x+z,q,a)\leq\frac{x+z}{\varphi(q)}+0.19  \frac{x+z}{\log (x+z)},\nonumber
	\end{equation}	
	and	
	\begin{equation}\label{eq3377D.150j}
		\frac{x}{\varphi(q)}-0.19  \frac{x}{\log x}\leq\psi(x,q,a)\leq\frac{x}{\varphi(q)}+0.19  \frac{x}{\log x},
	\end{equation}		
see \hyperlink{thm3377D.100}{Theorem} \ref{thm3377D.100} (i), yields 	
	\begin{eqnarray}\label{eq3377D.150r}
		&& \left(  \frac{x+z}{\varphi(q)} -(0.19)\frac{x+z}{\log (x+z)}\right)-\left( 	\frac{x}{\varphi(q)}-	(0.19)\frac{x}{\log x}\right)\nonumber\\[.3cm]	&\leq&\psi(x+z,q,a)-\psi(x,q,a)\\[.3cm]
		&\leq&\left(  \frac{x+z}{\varphi(q)} +(0.19)\frac{x+z}{\log (x+z)}\right)-\left( 	\frac{x}{\varphi(q)}+	(0.19)\frac{x}{\log x}\right)    \nonumber.
	\end{eqnarray}	
Rearranging and separating \eqref{eq3377D.150r} yield the claimed inequalities. Lastly, observe that the sums and differences in \eqref{eq3377D.150h} are well defined and nontrivial nonnegative real numbers for all $x^{\varepsilon}\leq z\leq x$ as $x\to\infty$.
\end{proof}

The smallest short interval $[x,x+z]$ possible is a topic of a long line of research, there are many conjectures and partial results. According to one of the earliest result, attributed to Westzynthius, almost every short interval $[x,x+z]$ with
\begin{equation}\label{eq3377D.150s}
z\ll \frac{\log x\log\log \log x}{\log \log\log \log x}
\end{equation}
is primefree ---  short intervals containing twin primes and prime $k$-tuples, where $k\geq2$ are small integers, are the exceptions. In contrast, almost every short interval $[x,x+z]$ with
\begin{equation}\label{eq3377D.150t}
	z\gg (\log x)^{1+\varepsilon},
\end{equation}
and $\varepsilon>0$, contains primes or is conjectured to contain primes, confer \cite{SA1943}, \cite{MS2007}, \cite{GL2023}, \cite{MT2024}, et alia for early results, surveys of the literature, numerical study and other recent developments on this topic. 

\section{Legendre Conjecture in Arithmetic Progressions }\label{S3377V}\hypertarget{S3377V}
The key to estimating the number of primes between two squares is the \textit{symmetric} explicit bounds of the prime counting function. Very recently proved symmetric explicit estimates are stated in \hyperlink{thm3377D.100}{Theorem} \ref{thm3377D.100}. The result in \hyperlink{thm3377D.050}{Theorem} \ref{thm3377D.050} is a simple application of this result. 

\begin{thm} \label{thm3377D.050}\hypertarget{thm3377D.050} Let $1\leq a<q\leq 10^5$, with $\gcd(a,q)=1$ be a pair of integers and let $n\geq 1$ be a large integer. Then, there exists a prime number $p\equiv a\bmod q$ in the range $n^2< p<(n+1)^2$. Furthermore, for a large number $x$, the number of weighted primes over the short interval $[x^2,(x+1)^2]$ has the lower bound
	$$	\psi((x+1)^2,q,a)-\psi(x^2,q,a)\geq \frac{2x}{\varphi(q)}+O\left( \frac{x}{\log x}\right).$$
\end{thm}

\begin{proof}[\textbf{Proof}] Let $X+Z=(x+1)^2$. By \hyperlink{lem3377D.150A}{Lemma} \ref{lem3377D.150A} it follows that there is an effective lower bound for the number of weighted primes over the short interval $[x^2,(x+1)^2]$ can be computed as follows. 
\begin{eqnarray}\label{eq3377D.200u}
			D(x)&=&	\psi((x+1)^2,q,a)-\psi(x^2,q,a)\\[.3cm]
&\geq&\left( \frac{(x+1)^2}{\varphi(q)}-	\frac{x^2}{\varphi(q)}   \right)               -0.19\left(	\frac{(x+1)^2}{\log (x+1)^2}
-\frac{x^2}{\log x^2} \right)\nonumber\\[.3cm]		
			&=&\frac{2x+1}{\varphi(q)}+E(x,q,a).\nonumber
			\end{eqnarray}
			
The error term $E(x,q,a)$ is estimated in \hyperlink{lem3377D.200S}{Lemma} \ref{lem3377D.200S}. Summing these estimates produces the lower bound
\begin{eqnarray}\label{eq3377D.200d}
\psi((x+1)^2,q,a)-\psi(x^2,q,a)
&\geq &	\frac{2x+1}{\varphi(q)}-0.19 \left( \frac{2x+1}{2\log x}+O\left( \frac{x}{(\log x)^2}\right)\right)\nonumber\\[.3cm]
&\geq &	\frac{2x}{\varphi(q)}+O\left( \frac{x}{\log x}\right)	\nonumber
\end{eqnarray}
for all large $x$ and small parameter $q=o(\log x)$. In particular, this is valid for $1\leq a<q\leq 10^5$, with $\gcd(a,q)=1$. Therefore, the quantity $D(x)$ is unbounded as $x\to\infty$.
\end{proof}


\begin{lem} \label{lem3377D.200S} \hypertarget{lem3377D.200S}If $x$ is a large number then 
	$$|E(x,q,a)|=0.19\left(	\frac{(x+1)^2}{\log (x+1)^2}
	-\frac{x^2}{\log x^2} \right) =0.19 \left( \frac{2x+1}{2\log x}+O\left( \frac{x}{(\log x)^2}\right)\right) .$$
\end{lem}	

\begin{proof}[\textbf{Proof}]			
	Expanding and simplifying the reciprocal of the first denominator yield
	\begin{eqnarray}\label{eq3377D.200k}
		\frac{1}{\log (x+1)^2}&=&\frac{1}{2[\log x+\log(1+1/x)]}\\[.3cm]
	&=&\frac{1}{2\log x}\cdot \frac{1}{\displaystyle 1+\frac{\log(1+1/x)}{\log x}}\nonumber\\[.3cm]
		&=&\frac{1}{2\log x}\cdot \frac{1}{\displaystyle 1+O\left( \frac{1}{x\log x}\right)}\nonumber\\[.3cm]
		&=&\frac{1}{2\log x}\left( 1+O\left( \frac{1}{x\log x}\right) \right) \nonumber.
	\end{eqnarray}
The second line in \eqref{eq3377D.200k} uses $\log(1+z)=O(z)$ for $|z|<1$
and the third line uses $1/(1+z)=1+O(z)$ for $|z|<1$, to complete the simplification. Replacing the new denominator \eqref{eq3377D.200k} back yields
	\begin{eqnarray}\label{eq3377D.200m}
		\frac{(x+1)^2}{\log (x+1)^2}
		-\frac{x^2}{\log x^2}
		&=&\frac{(x+1)^2}{2\log x}\left( 1+O\left( \frac{1}{x\log x}\right) \right) -\frac{x^2}{2\log x}\\[.3cm]
		&=&\frac{2x+1}{2\log x}+O\left( \frac{x^2+2x+1}{x(\log x)^2}\right) \nonumber\\[.3cm]
		&=&\frac{2x+1}{2\log x}+O\left( \frac{x}{(\log x)^2}\right)\nonumber .
	\end{eqnarray}
\end{proof}

\begin{thm} \label{thm3377U.050}\hypertarget{thm3377U.050} Let $1\leq a<q\leq 10^5$, with $\gcd(a,q)=1$ be a pair of integers and let $n\geq 1$ be a large integer. Then, there exists a prime number $p\equiv a\bmod q$ in the range $n^2< p<(n+1)^2$. Furthermore, for a large number $x$, the number of weighted primes over the short interval $[x^2,(x+1)^2]$ has the upper bound
	$$	\psi((x+1)^2,q,a)-\psi(x^2,q,a)\leq \frac{2x+1}{\varphi(q)}+O\left( \frac{x}{\log x}\right).$$
\end{thm}

\begin{proof}[\textbf{Proof}] The same routines as in the previous result mutatis mutandis.
\end{proof}

Merging the result in \hyperlink{thm3377D.050}{Theorem} \ref{thm3377D.050} and \hyperlink{thm3377U.050}{Theorem} \ref{thm3377U.050} leads to the inequalities
\begin{equation}
\frac{2x}{\varphi(q)}+O\left( \frac{x}{\log x}\right)\leq 	\psi((x+1)^2,q,a)-\psi(x^2,q,a)\leq \frac{2x+1}{\varphi(q)}+O\left( \frac{x}{\log x}\right).
\end{equation}

There are many numerical data on the Legendre conjecture, the most recent appears in \cite{SW2024}. 

\section{Primes Between Two Cubic Integers and Beyond }\label{S3377Y}  

A superficial inspection of the proof for primes between two squares given in \hyperlink{S3377V}{Section} \ref{S3377V} seems to work as well for primes between two cubic integers. Nevertheless, there is another method. The maximal gap $p_{n+1}-p_n\ll x^{5/8}$ between consecutive primes $p_n\leq x$, proved in \cite{IA1937}, implies the existence of primes between cubes, the most recent proof is in \cite{DA2016}, other cases are considered in \cite[Section 5]{CJ2023}. \\

An easy proof based on the maximal gap $p_{n+1}-p_n\ll x^{.525}$ between consecutive primes $p_n\leq x$ proved in \cite{BP2001} is assembled below.

\begin{thm} \label{thm3377Y.300}\hypertarget{thm3377Y.300} Let $1\leq a<q\leq 10^5$, with $\gcd(a,q)=1$ be a pair of integers and let $n\geq 1$ be a large integer. Then, there exists a prime number $p\equiv a\bmod q$ in the range $n^v< p<(n+1)^v$ for all $v\geq3$. 
	
\end{thm}	

\begin{proof}[\textbf{Proof}] By Theorem 1 in \cite{BP2001}, the interval $[x^v,x^v+x^{0.525v}]$ contains primes. Now observe that the binomial expansion of the $(x+1)^v=x^v+(v(v-1)/2)x^{v-1}+\cdots+1$, so there is a proper set inclusion of intervals
\begin{equation}
[x^v,x^v+x^{0.525v}]	\subset[x^v,x^v+cx^{v-1}]	\subset [x^v,(x+1)^v]
\end{equation}	
for any $v\geq3$ such that $0.525v< v-1$, where $c\geq v(v-1)\geq6$ is a small constant. The claim follows from these information.
\end{proof}



\end{document}